# Adapting SAM via Cross-Entropy Masking for Class Imbalance in Remote Sensing Change Detection

### v1

@@ Abstract
Foundational models have achieved significant success in diverse domains of computer vision. They learn general representations that are easily transferable to tasks not seen during training. One such foundational model is Segment anything model (SAM), which can accurately segment objects in images. We propose adapting the SAM encoder via fine-tuning for remote sensing change detection (RSCD) along with spatial-temporal feature enhancement (STFE) and multi-scale decoder fusion (MSDF) to detect changes robustly at multiple scales. Additionally, we propose a novel cross-entropy masking (CEM) loss to handle high class imbalance in change detection datasets. Our method outperforms state-of-the-art (SOTA) methods on four change detection datasets, Levir-CD, WHU-CD, CLCD, and S2Looking. We achieved 2.5% F1-score improvement on a large complex S2Looking dataset. The code is available at: https://github.com/humza909/SAM-CEM-CD
Index Terms:
Change Detection, Segment Anything Model, Vision Foundation Models, Remote Sensing
I Introduction
@@
Deep learning methods for change detection often employ the pre-trained weights of models trained on related vision tasks and domains to enhance performance. For example, the pre-trained ResNet backbone [8] was used to train bi-temporal image transformers for change detection (BIT) in [2], the pre-trained VGG [9] backbone was used in change-prior based change detection (CGNET) [6], and data-efficient transformers (deit) [10] were used in Changevit [3]. Although these backbone models are trained for different domains, their pre-learned features fine-tune well for remote sensing change detection, yielding significant performance gains compared to training these models from scratch [2, 6, 3]. Recently, the performance gains achieved by foundational models have attracted researchers to adapt them for change detection [11, 12]; for example, remote-clip’s [13] encoder was used in conjunction with change detection architectures [1, 2] to improve change detection performance.
@@
Figure 2: The architectural diagram of SAM-CEM-CD, X6 in classification head represents six repetitions of the residual module.
Most of the pre-trained models described above are trained for non-dense prediction tasks, such as image classification as opposed to image segmentation. Pixel-wise change detection is a dense segmentation task and the features extracted by encoders trained for similar dense spatial tasks are expected to be more effective for change detection. To this end, the Segment Anything Model (SAM) [14, 15] a foundational model for image segmentation was adapted for change detection in [11]. This work freezes the FastSAM (a smaller and faster variant of SAM) [15] encoder and trains an adapter and decoder from scratch, processing bi-temporal images sequentially and fusing them before classification.
This paper presents a new architecture for change detection that also utilizes pre-trained FastSAM. We extract multi-scale features from the feature pyramid network in FastSAM [15] for bi-temporal images. We employ a spatial-temporal feature enhancement (STFE) branch to learn cross-temporal relations at each scale. The output of the STFE branch is forwarded to a UNet-style decoder [16] followed by multi-scale decoder fusion (MSDF) and a residual head for change map classification. As opposed to [11], the encoder in our architecture uses the foundational model’s encoder connected in shared encoder/twin-style. Our use of fine-tuning on the FastSAM encoder outperforms change detection frameworks based on encoders using non-dense prediction models in [6, 2], as empirically evident in our experiments.
RSCD is a highly class imbalance problem, where pixels of change are significantly less frequent than non-change pixels. This class imbalance introduces an inherent bias in the learning process in which models emphasize improving the accuracy of non-change areas over areas of change. To address this, we propose a novel way of masking non-change pixels during the loss calculation, called cross-entropy masking (CEM), interchangeably cross-entropy dropout, shown in Figure 1.
@@
  • Proposing a SAM-based change detection architecture that learns cross-temporal multi-scale feature relations through STFE and MSDF for accurate change map classification.
  • We propose incorporating an encoder from a pre-trained model for dense image segmentation and fine-tuning it for change detection instead of using non-dense prediction models.
  • We propose a novel way to deal with class imbalance in change detection by masking or dropping non-change pixels in the ground-truth, naming it cross-entropy masking (CEM).
The rest of the paper is organized as follows: Section II discusses the proposed method, Section III discusses experimentation and results, followed by the conclusion in Section IV.
II Proposed Method
II-A Siamese FastSAM Encoder
@@
Given a set of two remote sensing images $I_{1}$ and $I_{2}$ for change detection, we process the images through the encoder. We extract four feature maps $F_{1}$, $F_{2}$, $F_{3}$, and $F_{4}$ from the encoder at four scales $1/4$, $1/8$, $1/16$, and $1/32$. The scale $1/4$ is extracted from the CNN backbone of FastSAM, where the remaining feature maps are taken from the feature pyramid network.
II-B Spatial-Temporal Feature Enhancement
We concatenate spatiotemporal representations and process them through transformations to learn cross-temporal relations and differences in feature maps, shown in Figure 2.
@@ II-B Spatial-Temporal Feature Enhancement
$F_{STFE}=SiLU(BN(Conv2D_{3x3}(F_{i}^{pre},F_{i}^{post})))$ (1)
@@
where $F_{STFE}$ are the enhanced feature maps, $F_{i}^{pre}$ and $F_{i}^{post}$ are pre-change and post-change images, respectively, $Conv2D$ is 2D convolution, $BN$ is batch normalization, and $SiLU$ is sigmoid linear unit activation function. The STFE module is replicated four times for feature maps at each scale extracted from the encoder. The enhanced representations are forwarded to the decoder for change detection.
T1 T2 GT SAM-CD SAM-CEM-CD (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Comparison of change detection results. T1 and T2 are the input images at different timestamps, GT is the ground truth, and SAM-CD [11] and SAM-CEM-CD (ours) are predictions from models. Red pixels are FPs and blue pixels are FNs.
II-C Unet Decoder
Multi-scale change features from the STFE module are used to reconstruct the change map in the UNet-style decoder [16]. The feature map resolution in the decoder gradually increases with skip connections, concatenating features from the STFE with the decoder output. The decoder block is written as:
$F_{up}$ $=ConvTrans2D_{2x2}(F_{STFE})$ (2)
$F_{out}$ $=SiLU(BN(Conv2D_{3x3}(F_{up},F_{dec}))$
@@ II-C Unet Decoder
In equation 2, the feature map from the difference branch (alternatively, STFE) $F_{STFE}$ is upsampled $F_{up}$ with transpose convolution $ConvTrans2D_{2x2}$ of kernel size $2\times 2$ to match the spatial resolution of the decoder output. The $F_{up}$ is concatenated with the previous output of the decoder to process through the next decoder block, providing contextual information on changes in the bitemporal images. For architectural consistency, we ensure that the decoder block has the convolution, batch normalization, and activation function in the same order as in the FastSAM encoder and the STFE module as given in equation 1.
II-D Multi-Scale Decoder Fusion and Change Map
@@
The early layers of the decoder contain rich global semantic information, whereas the last layers have local change structural detail. The decoder processes this information at multi-spatial resolutions, taking into account objects and changes in the temporal images at different scales. We collect every decoder block output, fuse it, and forward it to the classification head for change detection. Our classification head contains residual blocks followed by a convolution layer for the change map, as depicted in Figure 2.
II-E Cross-Entropy Masking for Class Imbalance
@@
The bitemporal change detection datasets are heavily imbalanced in change and unchanged regions. For instance, Levir-CD and WHU-CD have 95% unchanged pixels and only 5% change pixels. Because of this imbalance, the cross-entropy loss in change detection is highly biased towards accurately classifying unchanged pixels instead of change pixels. To ensure that the model emphasizes change pixels as equally as unchanged pixels, we randomly dropout the unchanged pixels from the binary cross-entropy loss. For this, we compute the per-element binary cross-entropy (BCE) as mentioned:
@@
TABLE I: Quantitative results of different RSCD methods on the Levir-CD and WHU-CD datasets, results are in %. The best results are highlighted.
@@
SAM-CEM-CD (ours) 96.12 95.84 99.23 95.98 92.51 97.45 97.19 99.54 97.23 94.73
III Experimentation and Results
TABLE II: Quantitative results of different RSCD methods on the CLCD dataset, results are in %.
@@ III Experimentation and Results
SAM-CEM-CD (ours) 88.71 86.89 96.71 87.76 79.77
TABLE III: Quantitative results of different RSCD methods on the S2Looking dataset for change class, results are in %.
@@
SAM-CEM-CD (ours) 73.20 63.02 67.73 51.19
TABLE IV: Ablation study using different modules in our architecture on the Levir-CD dataset, results are in %. STFE stands for spatio-temporal feature enhancement, D stands for Unet-decoder, MSDF stands for multi-scale decoder fusion, and CEM stands for cross-entropy masking.

### v2

# Multi-Receptive Field Ensemble with Cross-Entropy Masking for Class Imbalance in Remote Sensing Change Detection

@@ Abstract
Remote sensing change detection (RSCD) is a complex task, where changes often appear at different scales and orientations. Convolutional neural networks (CNNs) are good at capturing local spatial patterns but cannot model global semantics due to limited receptive fields. Alternatively, transformers can model long-range dependencies but are data hungry, and RSCD datasets are not large enough to train these models effectively. To tackle this, this paper presents a new architecture for RSCD which adapts a segment anything (SAM) vision foundation model and processes features from the SAM encoder through a multi-receptive field ensemble to capture local and global change patterns. We propose an ensemble of spatial-temporal feature enhancement (STFE) to capture cross-temporal relations, a decoder to reconstruct change patterns, and a multi-scale decoder fusion with attention (MSDFA) to fuse multi-scale decoder information and highlight key change patterns. Each branch in an ensemble operates on a separate receptive field to capture finer-to-coarser level details. Additionally, we propose a novel cross-entropy masking (CEM) loss to handle class-imbalance in RSCD datasets. Our work outperforms state-of-the-art (SOTA) methods on four change detection datasets, Levir-CD, WHU-CD, CLCD, and S2Looking. We achieved 2.97% F1-score improvement on a complex S2Looking dataset. The code is available at: https://github.com/humza909/SAM-ECEM
1 Introduction
@@
Figure 2: The architectural diagram of SAM-CEM-CD. The red, yellow, and green colored boxes represent an ensemble in spatio-temporal feature enhancement (STFE) and decoder, where each color operates on a separate receptive field from $1\times 1$, $3\times 3$, and $5\times 5$. The output of the red STFE boxes only connects with the red boxes in the decoder and so on.
Deep learning methods for change detection often employ pre-trained models as feature extractors. These pre-trained models are trained on large-scale vision datasets containing related tasks and domains. Because of this, the features extracted from these models are generalizable and transferable to other vision domains, such as change detection. ResNet [9] was used in bi-temporal image transformers for change detection (BIT) in [2], VGG [21] backbone was used in change-prior based change detection (CGNET) [8], and data-efficient transformers (deit) [22] were used in Changevit [28]. Following this, the FastSAM encoder, a segment anything model (SAM) variant, was used in SAM-CD [6]. It demonstrated better performance over past approaches by using an encoder trained on a dense prediction task for change detection, which itself is a dense prediction task, compared to an encoder trained on non-dense prediction tasks. These backbones fine-tune well for remote sensing change detection, where dataset sizes are comparatively smaller compared to other domains, yielding significant performance gains with respect to training these models from scratch [2, 8, 28].
Changes in remote sensing data occur on different scales with diverse object sizes and orientations. It is crucial for change detection models to capture finer and coarser details to accurately model bi-temporal changes. Many of the past approaches process multi-scale information through multi-resolution feature maps in the decoder [6, 1, 8, 17], however, the decoder runs through a single receptive field size, limiting it to capture details at a single granularity level.
Additionally, RSCD is a highly class-imbalanced problem, where pixels of change are significantly less frequent than non-change pixels. For example, Levir-CD [3] and WHU-CD [11] datasets have only 5% change pixels in the entire dataset. This class imbalance introduces an inherent bias in the learning process in which models emphasize improving the accuracy of non-change areas over areas of change. Loss functions such as focal loss [15], dice loss [10], and weighted cross-entropy [10] are suggested to handle class-imbalance. These loss functions adjust the importance of the majority and minority classes based on imbalance. However, the error for both classes propagates backwards, emphasizing more on the minority class that leads to overfitting and more false positives.
This paper presents a new RSCD architecture that addresses issues with class imbalance and local vs. global change information that are critical in change detection. We propose an architecture that uses an ensemble to learn features at different levels of granularity. Each branch in an ensemble operates on a different receptive field size to capture local and global details, which are combined for classification. We also propose a loss mechanism for class imbalance, which masks the loss calculation pixel-by-pixel for classes in the majority, as shown in Figure 1. Our architecture initially extracts multi-resolution features from a vision foundation model FastSAM [27] and learns cross-temporal relations using spatio-temporal feature enhancement (STFE), followed by a decoder ensemble to refine and upsample change patterns. The multi-scale decoder fusion with attention (MSDFA) fuses multi-scale information from the decoder and highlights important change regions for the classification head.
@@
  • We propose an ensemble of STFE, decoder, and MSDFA, where each branch operates on a range of receptive field sizes to capture local and global contextual changes at different scales.
  • We propose a novel cross-entropy masking (CEM) loss that randomly masks pixels for the majority class during loss calculation to reduce class imbalance for change detection, outperforming other class-imbalance losses.
  • We demonstrate the effectiveness of our architecture on four remote sensing change detection (RSCD) datasets, outperforming previous state-of-the-art (SOTA) methods in RSCD.
2 Related Work
Architectures in RSCD are single encoder-decoder [5], dual encoder single decoder [1, 8], or dual encoder-decoder [6] architectures, which are mainly divided into three categories: 1) early fusion, 2) middle fusion, and 3) late fusion. In early fusion [5], temporal images are concatenated before processing through a deep learning model. In middle fusion [5, 8, 17], features from both encoder branches are combined to learn cross-temporal relationships, whereas in late fusion [6] features are processed through encoder-decoder branches and merged together at the end for change map classifcation.
Temporal images were concatenated in early convolutional neural networks (CNNs) based change detection methods [5]. The input data was processed through a UNet-style encoder-decoder architecture containing skip connections with multi-scale information processing. In another siamese-style variant for change detection by the same work [5], features extracted at multi-resolutions are subtracted to highlight change features. However, simple concatenation of bi-temporal multi-resolution features in difference branch was shown to be better to learn change features [1, 7]. Instead of sparse connections, a densely connected UNet change detection architecture was proposed in [7]. Change priors were used to guide change detection in [8]. This work also uses pre-trained backbones such as VGG [21] instead of training feature extractors from scratch. To emphasize key change regions, channel attention, spatial attention, or other attention mechanisms were used in [14, 4, 20].
The self-attention in transformers can capture global contextual dependencies opposite to CNNs. Based on this, many change detection methods were proposed using transformers [28, 25, 17]. A combination of CNNs and transformers were proposed in [2, 26], extracting initial features from ResNet [9] followed by transformers-based encoder-decoder architecture. ChangeFormer [1] introduced a hierarchical transformer encoder to learn multi-resolution features similar to CNNs. However, transformers lack CNNs inductive bias and translational invariance, making them data hungry. Furthermore, the RSCD datasets [3, 11] are not large enough to train models from scratch.
Because of this, researchers have started adopting pre-trained vision foundational models for RSCD. Generalized features extracted from these models are used to train robust change detection models. In [13], features from the remote-clip model are concatenated with the encoders in ChangeFormer [1] and BIT [2] models for better generalization. Segment anything model (SAM) [12] CNN-based variant FastSAM [27] was used in [6]. Features from the FastSAM encoder are used to reconstruct bi-temporal images through a decoder followed by late fusion and classification.
Following this, our work employs the FastSAM vision foundation model for change detection. We use an ensemble of smaller and wider receptive fields to capture local and global structural changes in the data. Additionally, we propose a cross-entropy masking loss for a better generalization of the model.
3 Proposed Method
3.1 Siamese FastSAM Encoder
@@
Given a set of two remote sensing images $I_{1}$ and $I_{2}$ for change detection, we process the images through the encoder. We extract four feature maps $F_{1}$, $F_{2}$, $F_{3}$, and $F_{4}$ from the encoder at four scales $1/4$, $1/8$, $1/16$, and $1/32$.
3.2 Spatial-Temporal Feature Enhancement
We concatenate features extracted at each scale to learn cross-temporal relations and differences in feature maps. Change patterns within a single resolution feature map occur at different sizes; therefore, it is important to learn spatio-temporal representations at different contextual levels. Contrary to the work in [1, 6, 8, 7], our spatial-temporal feature enhancement (STFE) module processes the input using an ensemble of convolutional receptive fields of sizes $1\times 1$, $3\times 3$, and $5\times 5$ to model changes at various scales. We show the effectiveness of using a multi-receptive field ensemble compared to a single receptive field in Table 5. The STFE module is replicated four times for different resolution feature maps, as shown in Figure 2.
@@ 3.2 Spatial-Temporal Feature Enhancement
$F_{i}^{STFE}=SiLU(BN(Conv2DWS_{k\times k}(F_{j}^{pre},F_{j}^{post})))$ (1)
@@
where $F_{i}^{STFE}$ are the enhanced feature maps, $F_{j}^{pre}$ and $F_{j}^{post}$ are pre-change and post-change images, respectively, $Conv2DWS$ is 2D depthwise separable convolution, $BN$ is batch normalization, $SiLU$ is sigmoid linear unit activation function, $i$ is equal to 1, 3, or 5 representing the convolutional receptive field, and $j$ represents the feature map output from the encoder.
3.3 Decoder Ensemble
The decoder in our architecture is an ensemble of UNet-style decoders [18]. Each decoder in ensemble operates in one of the convolutional receptive fields of $1\times 1$, $3\times 3$, and $5\times 5$, and receives input from the $F_{i}^{STFE}$ output of the respective receptive field. This design of the model allows for emphasizing and reconstructing changes at different level of details, finer to coarse. The decoder block is written as:
$F_{up}$ $=ConvTrans2D_{2x2}(F_{i}^{STFE})$ (2)
$F_{out}$ $=SiLU(BN(Conv2D_{i\times i}(F_{up},F_{dec}))$
@@ 3.3 Decoder Ensemble
In equation 2, the feature map from the STFE branch $F_{i}^{STFE}$ is upsampled $F_{up}$ with transposed convolution $ConvTrans2D_{2x2}$ of kernel size $2\times 2$ to match the spatial resolution of the decoder output. The $F_{up}$ is concatenated with the previous output of the decoder to process through the next decoder block, providing contextual information on changes in the bitemporal images. We empirically found that using standard 2D convolution in decoder performs better than depthwise separable convolution, which allows cross-channel interaction for better reconstruction.
3.4 Multi-Scale Decoder Fusion with Attention
@@
The early layers of the decoder contain rich global semantic information, whereas the last layers have local change structural detail. The decoder processes this information at multi-spatial resolutions, taking into account objects and changes in the temporal images at different scales.
Multi-scale decoder fusion with attention (MSDFA) concatenates the output of the decoder blocks, shown in equation 3.
$x_{i}=D_{2,i}\oplus upsample(D_{1,i})\oplus upsample(D_{0,i})$ (3)
Upsample is a bilinear interpolation operation to match the feature resolution of the early decoder blocks to the last decoder block. The $i$ represents the decoder ensemble number based on the receptive field sizes 1, 3, or 5. Before attention, we inject local context using depthwise separable convolution on $x_{i}$ with a receptive field size same as the decoder. Afterwards, we apply channel attention followed by the spatial attention as given in equations 4 and 5, respectively.
$z_{i}=x_{i}\otimes\sigma(Conv1D_{1\times k}(GAP(x_{i})))$ (4)
$y_{i}=z_{i}\otimes\sigma(Conv2D_{k\times k}(Avg(z_{i})\oplus Max(z_{i})))$ (5)
In equation 34 and 5, $\oplus$ represents concatenation, $\otimes$ is elementwise multiplication, $\sigma$ is sigmoid activation function, $GAP$ is global average pooling, $Avg$ and $Max$ are channel-wise average and maximums. To mention here, $k$ is equal to 3, 5, and 7 when $i$ is equal to 1, 3, and 5, respectively. We select a larger odd-numbered receptive field in attention modules compared to the original receptive field in the ensemble. This allows attending on a wider context to highlight key change areas and performs better than keeping a single kernel size for all $i$’s.
The output $y_{i}$ is summed with the input $x_{i}$, making it a residual connection and written as $y_{i}=y_{i}+x_{i}$. Our attention module is similar to the convolution block attention module (CBAM) [23], except our channel attention uses only GAP and convolution instead of a multi-layer perceptron layer (MLP). Additionally, we have a separate MSDFA for each decoder in the decoder ensemble. The final output from the MSDFA is given as:
$y=y_{1}\oplus y_{3}\oplus y_{5}$ (6)
3.5 Classification Head
The last layer in our architecture is the classification head to generate change maps. It comprises six residual layers followed by a convolutional classification layer. The input to the classification head is the concatenated output from the MSDFA layer.
3.6 Cross-Entropy Masking for Class Imbalance
@@
The bitemporal change detection datasets are heavily imbalanced in change and unchanged regions. For instance, Levir-CD [3] and WHU-CD [11] have 95% unchanged pixels and only 5% change pixels. Because of this imbalance, the cross-entropy loss in change detection is highly biased towards accurately classifying unchanged pixels instead of change pixels. To ensure that the model emphasizes change pixels as equally as unchanged pixels, we randomly dropout the unchanged pixels from the binary cross-entropy loss. For this, we compute the per-element binary cross-entropy (BCE) as mentioned:
@@
Table 1: Quantitative results of different RSCD methods on the Levir-CD and WHU-CD datasets, results are in %. The best results are highlighted.
@@
SAM-ECEM (ours) 96.27 95.72 99.23 96.01 92.56 97.82 96.69 99.54 97.25 94.76
4 Experimentation and Results
Table 2: Quantitative results of different RSCD methods on the CLCD dataset, results are in %.
@@ 4 Experimentation and Results
SAM-ECEM (ours) 88.43 87.53 96.73 88.04 80.14
Table 3: Quantitative results of different RSCD methods on the S2Looking dataset for change class, results are in %.
@@
SAM-ECEM (ours) 74.59 62.65 68.10 51.63
4.1 Implementation Details
We used PyTorch for model training. We trained the model for 35 epochs with three additional warm-up epochs using a stochastic gradient descent (SGD) optimizer. We use the FastSAM encoder and fine-tune its pre-trained weights. The learning rate (lr) is 0.01 for Levir-CD and CLCD, and 0.001 for WHU-CD and S2Looking. We decay the learning rate in each iteration $lr*(1-iterations/50)^{decay}$, where decay is 2.0 for Levir-CD and CLCD, and 3.0 for WHU-CD and S2Looking. Image flipping and random cropping augmentations are applied during training and images are flipped 4 times for inference to produce stable results. More details can be seen in our GitHub repo: https://github.com/humza909/SAM-ECEM.
4.2 Change Detection Datasets
Levir-CD [3]: This dataset contains 637 image pairs with 31,333 annotated change objects. The data is collected from Google Earth for six Texas cities between 2012 and 2016. It has 445 training, 64 validation, and 128 testing images.
WHU-CD [11]: This is an aerial imaging dataset captured over Christchurch, New Zealand, after the 2011 earthquake, between 2012-2016. We used the dataset split of 6096 training, 762 validation, and testing images provided by [6].
CLCD [24]: The cropland change detection dataset is captured in Guangdong, China between 2017 to 2019. It consists of 600 image pairs collected by the Gaofen-2 satellite with a split of 320 training, 120 validation and testing image pairs.
S2Looking [19]: This is a large-scale change detection dataset with 5000 image pairs collected from satellites such as Gaofen, SuperView, and BeiJing-2. The dataset presents various challenges, including misalignment and varying poses.
4.3 Evaluation Metrics
To evaluate the performance of our model, we employ widely used metrics for change detection, including precision (Prec), recall (Rec), $F_{1}$ score, overall accuracy (OA), and intersection over union (IoU). Tables 1 and 2 contain the average of F1, represented as mF1, and other metrics for the background and foreground classes, whereas in Table 3, the results are reported only for the foreground class.
T1 T2 GT CF BIT CGNet SAM-CD Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Comparison of change detection results. T1 and T2 are the input images at different timestamps, GT is the ground truth, and CF [1], BIT [2], CGNet [8], SAM-CD [6], and SAM-ECEM (ours) are predictions from models. Red pixels are FPs and blue pixels are FNs.
4.4 Quantitative Results
Our method, SAM-ECEM, achieves superior performance than other SOTA methods in the literature. For the correctness of the results, we follow the same evaluation strategy as in [6]. The results in Tables 12, and 3 are also collected from [6]. Table 1 shows an increase of 0.51% and 0.88% in mF1 and mIoU, respectively, for the Levir-CD dataset. Similarly, we achieve 97.25% mF1 and 94.76% mIoU for the WHU-CD dataset, compared to the second-best CGNET [8], which yields 97.07% mF1 and 94.44% mIoU. We recalculated the SAM-CD results for WHU-CD as given in Table 1. Table 2 shows an approximately 1.5% improvement in mF1 and mIoU for CLCD against other methods, whereas for S2Looking, we achieve 2.97% higher F1 and 3.34% higher IoU performance than the SOTA methods in Table 3.
4.5 Visualization
We visually compare the change maps of ChangeFormer [1], BIT [2], CGNeT [8], SAM-CD [6] and SAM-ECEM (ours) in Figure 3. The blue predictions are FNs and the red predictions are FPs. We can see the reduced FNs and FPs by SAM-ECEM in the figure, which shows the significance of our method over other SOTA methods.
4.6 Ablation Study
Table 4: Ablation study using different modules in our architecture on the Levir-CD dataset, results are in %. STFE stands for spatio-temporal feature enhancement, D stands for Unet-decoder, MSDFA stands for multi-scale decoder fusion with attention, and CEM stands for cross-entropy masking.